Adult Americans (18 and older) were asked whether they use social media regularly. The following table is based on the results of the survey.

What is the probability the person used social media given that they were between the ages 35 and 44? Round your answer to three decimal places.

18-34

35-44

45-54

55+

Use social media

117

89

83

49

Do not use social media

33

36

57

66

Group of answer choices

.638 ,

1.404

.712

1.568
Answer #1

From the given information,

By using calculator,

The required correct answer is,

P(used social media | 35-44)= 89/(89+36)= 89/125 = 0.712

i.e., Third option is correct.

Please help me out by just rating the answer.

Thank you so much.